Circular No. 6062

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

JUPITER AND COMET 1993e S. Sukumar, Dominion Radio Astrophysical Observatory, Penticton,

reports the preliminary results of radio monitoring of Jupiter during July 15–21 at a frequency of 1420 MHz, using the synthesis radio telescope: “Jupiter was continuously tracked for ∼ 6.6 hr each day centered around ∼ 2^{h}30^{m} UT; the calibrated data were mapped and the total intensity of Jupiter was measured on each day within a rms noise limit of ∼ 60 mJy. The apparent total flux density of Jupiter, uncorrected for distance, was ∼ 3.5 Jy on July 15 and 16 but increased steadily after July 17, reaching a peak level of 4.05 Jy on July 20. A slow decrease to a level of 3.89 Jy on July 21 was measured. These measurements are consistent with similar variation observed with the ATNF compact array at 1430 MHz (Kesteven et al., IAUC 6040); in conjunction with the increase of 50 percent observed at 843 MHz with the MOST (Hunstead and Campbell-Wilson, IAUC 6038), this might indicate an enhancement in the nonthermal radio emission from Jupiter during the infall of the fragments of comet 1993e.”

X-RAY NOVA IN SCORPIUS R. Hunstead, University of Sydney, on behalf of D. Campbell-Wilson

(University of Sydney and Molonglo Observatory Synthesis Telescope), D. McKay (Australia Telescope National Facility = ATNF), and J. Lovell (University of Tasmania), reports preliminary results from the GRO J1655–40 flux-monitoring program during Aug. 15–23 UT at frequencies 0.843–9.2 GHz: “At 843 MHz, following a brief maximum at Aug. 15.2–15.3 UT (5.5 Jy), the flux density fell to a minimum around Aug. 16.1 (3.9 Jy), rising again to a peak around Aug. 17.2 (7 Jy) that persisted for 1.3 days. Since then, the 843-MHz flux density has decayed steadily, reaching 2.6 Jy on Aug. 22.6. The data at higher frequencies (available after Aug. 15.3) have followed similar trends, with a spectral index of −0.6 at the Aug. 17 maximum, flattening slightly to −0.4 at Aug. 22.6. Apart from the period of rapid rise (Aug. 16–17), the spectrum has remained a power-law over the full frequency range.”

D. McKay and M. Kesteven, ATNF, report: “H I observations of GRO J1655–40 made with the AT Compact Array show solid absorption in the velocity range +10 to −30 km/s, with a further isolated weak feature at −50 km/s. The balance of probabilities is that the distance is around 3.5 kpc, unless the −50 km/s feature is due to an atypical cloud.”

1994 August 26 Daniel W. E. Green